# Kernel PCA for Out-of-Distribution Detection

### v1

@@ 3 Methodology
3.1 Preliminary: PCA for OoD Detection
$\boldsymbol{\Sigma}=\sum_{i=1}^{N_{\rm tr}}\left(\boldsymbol{z}_{i}-\boldsymbol{\mu}\right)\left(\boldsymbol{z}_{i}-\boldsymbol{\mu}\right)^{\top},\ \boldsymbol{\mu}=\frac{1}{N_{\rm tr}}\sum_{i=1}^{N_{\rm tr}}\boldsymbol{z}_{i}.$ (3)
Through the eigendecomposition $\boldsymbol{\Sigma}=\boldsymbol{U}\boldsymbol{\Lambda}\boldsymbol{U}^{\top}$, the dimensionality reduction matrix $\boldsymbol{U}_{q}\in\mathbb{R}^{m\times q}$ is obtained by taking the first $q$ columns of the eigenvector matrix $\boldsymbol{U}$ w.r.t. the top-$q$ eigenvalues.
In inference, given a new sample $\boldsymbol{\hat{x}}\in\mathbb{R}^{d}$ and its feature $\boldsymbol{\hat{z}}\in\mathbb{R}^{m}$ from DNN $f$, the reconstruction error is computed:
By projecting centralized $(\boldsymbol{\hat{z}}-\boldsymbol{\mu})$ to the $\boldsymbol{U}_{q}$-subspace and projecting back, we can obtain the reconstructed features $\boldsymbol{U}_{q}\boldsymbol{U}_{q}^{\top}(\boldsymbol{\hat{z}}-\boldsymbol{\mu})$ and the reconstruction error $e(\boldsymbol{\hat{x}})$, which then can be set as the OoD detection score: $S(\boldsymbol{\hat{x}})=-e(\boldsymbol{\hat{x}})$. An ideal case is that $\boldsymbol{U}_{q}$ contains informative principal components of InD data and causes projections of OoD data far away from that of InD data, leading to separable reconstruction errors between OoD and InD data.
3.2 Kernel PCA for OoD Detection
As empirically observed in [8], the aforementioned PCA reconstruction error in the $\boldsymbol{z}$-space is not an effective score in detecting OoD data from InD data, which results from the weak ability of PCA in handling linearly inseparable data distributions. To address this issue, we propose to explore the non-linearity in $\boldsymbol{z}$-space via kernel PCA. Then, through a kernel perspective on an existing KNN detector [7], we put forward two feature mappings $\Phi$ w.r.t. two kernels: a cosine kernel (Sec.3.2.1) and a cosine-Gaussian kernel (Sec.3.2.2). The devised feature mappings $\Phi$ well characterize the non-linear patterns in $\boldsymbol{z}$-space of InD and OoD data, and thus benefit PCA in the mapped $\Phi(\boldsymbol{z})$-space for extracting an informative principal subspace and producing distinguishable reconstruction errors, as illustrated in Fig.1.
3.2.1 Cosine Kernel
In the KNN detector [7], the nearest neighbor searching is executed on the $\ell_{2}$-normalized penultimate features, i.e., $\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}$. In inference, given a new sample $\boldsymbol{\hat{x}}$, it operates the $\ell_{2}$ normalization by $\frac{\boldsymbol{\hat{z}}}{\|\boldsymbol{\hat{z}}\|_{2}}$ and the negative of its ($k$-th) shortest $\ell_{2}$ distance to $\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}$ of training samples is set as the detection score:
@@ 3.2.1 Cosine Kernel
The ablations in KNN demonstrate the indispensable significance of the $\ell_{2}$-normalization: the nearest neighbor searching directly on $\boldsymbol{z}$ shows a notably drop in detection performance. The critical role of the $\ell_{2}$-normalization in KNN attracts our attention in the sense of kernel. From a kernel perspective, the $\ell_{2}$-normalization is the non-linear feature mapping $\phi_{\rm cos}$ inducing the cosine kernel $k_{\rm cos}$:
$k_{\rm cos}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=\frac{\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{2}}{\|\boldsymbol{z}_{1}\|_{2}\cdot\|\boldsymbol{z}_{2}\|_{2}},\quad\phi_{\rm cos}(\boldsymbol{z})=\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}.$ (6)
@@
It indicates that a justified $\phi_{\rm cos}(\boldsymbol{z})$-space with non-linear mapping, instead of the original $\boldsymbol{z}$-space, contributes to the success of nearest neighbor searching in detecting OoD.
@@
Therefore, notice that the key of KPCA for OoD detection lies in a non-linear feature space, either through the kernel $k$ or the associated explicit feature mapping $\Phi$. Motivated by the KNN detector, we apply $\Phi(\cdot)\triangleq\phi_{\rm cos}(\cdot)$ as the feature mapping in KPCA to introduce non-linearity. Then, the PCA is executed on mapped features $\phi_{\rm cos}(\boldsymbol{z})$, following the procedures described in Sec.3.1. All the features $\boldsymbol{z}$ are now mapped to $\Phi(\boldsymbol{z})$ to formulate the covariance matrix $\boldsymbol{\Sigma}^{\Phi}$, for computing non-linear principal components with matrix $\boldsymbol{U}_{q}^{\Phi}$ and the corresponding reconstruction error $e^{\Phi}$. This detection scheme is dubbed as CoP (Cosine mapping followed by PCA), as in Algorithm 1.
@@
8:Calculating the covariance matrix: $\boldsymbol{\Sigma}^{\Phi}=\sum_{i=1}^{N_{\rm tr}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}{\Phi(\boldsymbol{z}_{i})}}-\boldsymbol{\mu}^{\Phi})({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}{\Phi(\boldsymbol{z}_{i})}}-\boldsymbol{\mu}^{\Phi})^{\top}$, $\boldsymbol{\mu}^{\Phi}=\frac{1}{N_{\rm tr}}\sum_{i=1}^{N_{\rm tr}}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}{\Phi(\boldsymbol{z}_{i})}$.
@@
10:Taking the first $q$ columns of $\boldsymbol{U}^{\Phi}$: $\boldsymbol{U}^{\Phi}_{q}=\boldsymbol{U}^{\Phi}\left[:,:q\right]$.
11:Dimensionality-reduction matrix $\boldsymbol{U}_{q}^{\Phi}$.
12:Given a new sample $\boldsymbol{\hat{x}}$ and its features $\boldsymbol{\hat{z}}$.
13:Calculating the reconstruction error: $e^{\Phi}(\boldsymbol{\hat{x}})=\left\|\boldsymbol{U}^{\Phi}_{q}\boldsymbol{U}_{q}^{\Phi\top}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}{\Phi(\boldsymbol{\hat{z}})}}-\boldsymbol{\mu}^{\Phi})-({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}{\Phi(\boldsymbol{\hat{z}})}}-\boldsymbol{\mu}^{\Phi})\right\|_{2}$.
14:Reconstruction error $e^{\Phi}$.
3.2.2 Cosine-Gaussian Kernel
The success of KNN, i.e., Eq.(5), suggests that the $\ell_{2}$ distance on $\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}$ is effective in distinguishing OoD data from InD data. In other words, the $\ell_{2}$ distance relation between samples in the $\phi_{\rm cos}$-space preserves useful information that benefits the separation of OoD data from InD data. This motivates us to introduce non-linear feature spaces that can retain the $\ell_{2}$ distance relation. Hence, we propose to introduce KPCA with non-linearity built upon $\phi_{\rm cos}(\boldsymbol{z})$, through which the useful $\ell_{2}$ distance in $\phi_{\rm cos}$-space can be preserved to further separate InD and OoD data.
In this regard, we deploy the shift-invariant Gaussian kernel:
$k_{\rm gau}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=e^{-\gamma\|\boldsymbol{z}_{1}-\boldsymbol{z}_{2}\|^{2}_{2}},$ (7)
where the $\ell_{2}$ distance information can be kept. The feature mapping associated with $k_{\rm gau}$ is infinite-dimensional, but it can be efficiently approximated through random Fourier features [27] (RFFs), i.e., $\phi_{\rm RFF}$ defined in Eq.(2). In this way, the inner product of two mapped samples $\phi_{\rm RFF}(\boldsymbol{z}_{1})^{\top}\phi_{\rm RFF}(\boldsymbol{z}_{2})$ provides the approximate Gaussian kernel, so that we can leverage the RFF feature mapping $\phi_{\rm RFF}$ to preserve the $\ell_{2}$ distance information in the Gaussian kernel, which has been verified effective for separating OoD data from InD data. We name it as cosine-Gaussian kernel, as the Gaussian kernel $k_{\rm gau}$ (or $\phi_{\rm RFF}$) is imposed on top of the cosine kernel $k_{\rm cos}$ (or $\phi_{\rm cos}$), further exploiting the $\ell_{2}$ distance relationships beyond the $\phi_{\rm cos}$-space for OoD detection. As we work with the explicit feature mapping, the non-linearity to $\boldsymbol{z}$ is achieved by $\Phi(\cdot)\triangleq\phi_{\rm RFF}(\phi_{\rm cos}(\cdot))$. With mappings $\phi_{\rm RFF}(\phi_{\rm cos}(\boldsymbol{z}))$, PCA is then executed to compute the reconstruction errors for OoD detection. This detection scheme is dubbed as CoRP (Cosine and RFFs mappings followed by PCA). Algorithm 1 illustrates the complete procedure of the proposed CoP and CoRP for OoD detection.
To warp up, we devise two effective feature mappings w.r.t. a cosine kernel and a cosine-Gaussian kernel to promote the separability of InD data and OoD data in non-linear feature spaces, inspired by effectiveness of the $\ell_{2}$ normalization and the $\ell_{2}$ distance from a kernel perspective on the KNN detector [7]. Our two feature mappings well characterize the non-linearity in penultimate features $\boldsymbol{z}$ of DNNs between InD and OoD data, enabling PCA to extract an informative subspace w.r.t. the mapped features through principal components and the reconstruction errors.
Notice that in our method the matrix $\boldsymbol{U}_{q}^{\Phi}$ is calculated on the InD training samples. Given any new test sample with the penultimate feature $\boldsymbol{z}$, in inference, we only need the feature mapping and reconstruction error of $\boldsymbol{\hat{z}}$ itself, leading to a constant time complexity $\mathcal{O}(1)$, which is significantly more efficient than the $\mathcal{O}(N_{\rm tr})$ in KNN. Numerical supports are provided in the following Sec.4.
@@ 3.2.2 Cosine-Gaussian Kernel
Table 1: The detection performance of different methods (ResNet50 trained on ImageNet-1K) on each OoD data set.
method OoD data sets
iNaturalist SUN Places Textures AVERAGE
@@
Without Supervised Contrastive Learning
MSP 54.99 87.74 70.83 80.86 73.99 79.76 68.00 79.61 66.95 81.99
ODIN 47.66 89.66 60.15 84.59 67.89 81.78 50.23 85.62 56.48 85.41
Energy 55.72 89.95 59.26 85.89 64.92 82.86 53.72 85.99 58.41 86.17
GODIN 61.91 85.40 60.83 85.60 63.70 83.81 77.85 73.27 66.07 82.02
Mahalanobis 97.00 52.65 98.50 42.41 98.40 41.79 55.80 85.01 87.43 55.47
KNN 59.00 86.47 68.82 80.72 76.28 75.76 11.77 97.07 53.97 85.01
@@
With Supervised Contrastive Learning
SSD 57.16 87.77 78.23 73.10 81.19 70.97 36.37 88.52 63.24 80.09
KNN 30.18 94.89 48.99 88.63 59.15 84.71 15.55 95.40 38.47 90.91
@@
4 Experiments
In experiments, our KPCA-based detectors, CoP and CoRP, are firstly compared with the KNN detector [7] in Sec.4.1. In Sec.4.2, CoP and CoRP are further compared with the regularized reconstruction error proposed in [8], and achieve SOTA OoD detection performance among various prevailing methods on multiple OoD data sets and network structures. A sensitivity analysis on the hyper-parameters of CoP and CoRP is presented in Sec.4.3 and ablation studies are provided in Sec.4.4.
The experiments are executed on the commonly-used small-scale CIFAR10 benchmark [31] and the large-scale ImageNet-1K benchmark [32].
  • For CIFAR10 as the in-distribution, the out-distribution data sets include SVHN [33], LSUN [34], iSUN [35], Textures [36] and Places365 [37].
  • For ImageNet-1K as the in-distribution, the out-distribution data sets consist of iNaturalist [38], SUN [39], Places [37] and Textures [36].
For the evaluation metrics, we employ the commonly-used (i) False Positive Rate of OoD samples with 95% true positive rate of InD samples (FPR95), and (ii) Area Under the Receiver Operating Characteristic curve (AUROC). The source code has been released [^1] https://github.com/fanghenshaometeor/ood-kernel-pca.

### v3

@@ 4 Methodology
As empirically observed in [8], the aforementioned PCA reconstruction error in the $\boldsymbol{z}$-space is not an effective score in detecting OoD data from InD data. We propose that the reason behind is possibly due to the linearly-inseparable features of InD and OoD data, as shown in Figure 1a. To address this issue, we propose to explore the non-linearity in $\boldsymbol{z}$-space via kernel PCA. Then, through a kernel perspective on an existing KNN detector [7], we put forward two efficacious kernels that well characterize the non-linear patterns in $\boldsymbol{z}$-space of InD and OoD data: a cosine kernel (Section 4.1) and a cosine-Gaussian kernel (Section 4.2). Particularly, we adopt two explicit feature mappings $\Phi$ induced from the two kernels, and execute PCA in the mapped $\Phi(\boldsymbol{z})$-space, which leads to an informative principal subspace and distinct reconstruction errors for efficacious OoD detection.
4.1 Cosine kernel
In the KNN detector [7], the nearest neighbor searching is executed on the $\ell_{2}$-normalized penultimate features, i.e., $\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}$. In inference, given a new sample $\boldsymbol{\hat{x}}$, its feature $\boldsymbol{\hat{z}}$ is firstly normalized as $\frac{\boldsymbol{\hat{z}}}{\|\boldsymbol{\hat{z}}\|_{2}}$, then the negative of its ($k$-th) shortest $\ell_{2}$ distance to the $\ell_{2}$-normalized features $\frac{\boldsymbol{z}_{i}}{\|\boldsymbol{z}_{i}\|_{2}}$ of training data is set as the detection score:
@@ 4.1 Cosine kernel
The ablations in KNN demonstrate the indispensable significance of the $\ell_{2}$-normalization: the nearest neighbor searching directly on $\boldsymbol{z}$ shows a notably drop in detection performance. The critical role of the $\ell_{2}$-normalization in KNN attracts our attention in the sense of kernel. From a kernel perspective, the $\ell_{2}$-normalization is exactly the non-linear feature mapping $\phi_{\rm cos}$ inducing the cosine kernel $k_{\rm cos}$:
$k_{\rm cos}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=\frac{\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{2}}{\|\boldsymbol{z}_{1}\|_{2}\cdot\|\boldsymbol{z}_{2}\|_{2}}=\phi_{\rm cos}(\boldsymbol{z}_{1})^{\top}\phi_{\rm cos}(\boldsymbol{z}_{2}),\quad\phi_{\rm cos}(\boldsymbol{z})=\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}.$ (6)
@@
It indicates that a justified $\phi_{\rm cos}(\boldsymbol{z})$-space with such non-linear mapping, instead of the original $\boldsymbol{z}$-space, contributes to the success of nearest neighbor searching in detecting OoD.
@@
Notice that the key of KPCA for OoD detection lies in an appropriate non-linear feature space that captures the non-linearity in InD and OoD features, either through the kernel $k$ or the associated explicit feature mapping $\Phi$. Motivated by the KNN detector, we apply $\Phi(\cdot)\triangleq\phi_{\rm cos}(\cdot)$ as the feature mapping in KPCA to introduce non-linearity. Then, PCA is executed on mapped features $\phi_{\rm cos}(\boldsymbol{z})$, following the procedures described in Section 3.1. All the features $\boldsymbol{z}$ are now mapped to $\Phi(\boldsymbol{z})$ to formulate the covariance matrix $\boldsymbol{\Sigma}^{\Phi}$, for computing non-linear principal components with matrix $\boldsymbol{U}^{\Phi}_{q}$ and the corresponding reconstruction error $e^{\Phi}$. This detection scheme is dubbed as CoP (Cosine mapping followed by PCA), as shown in Algorithm 1. An in-depth analysis on the effect of the normalization of the cosine kernel is left in Section C.1.
4.2 Cosine-Gaussian kernel
The success of KNN (Equation 5) suggests that the $\ell_{2}$ distance on $\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}$ is effective in distinguishing OoD data from InD data. In other words, the $\ell_{2}$ distance relation between samples in the $\phi_{\rm cos}$-space preserves useful information that benefits the separation of OoD data from InD data. This motivates us to seek non-linear feature spaces that can retain the $\ell_{2}$ distance relation. Hence, we propose to introduce KPCA with non-linearity built upon $\phi_{\rm cos}(\boldsymbol{z})$, through which the useful $\ell_{2}$ distance in $\phi_{\rm cos}$-space can be preserved to further separate InD and OoD data.
In this regard, we deploy the shift-invariant Gaussian kernel to keep the $\ell_{2}$ distance information:
$k_{\rm gau}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=e^{-\gamma\|\boldsymbol{z}_{1}-\boldsymbol{z}_{2}\|^{2}_{2}}.$ (7)
The feature mapping associated with $k_{\rm gau}$ is infinite-dimensional, but it can be efficiently approximated through random Fourier features (RFFs, [13]), i.e., $\phi_{\rm RFF}$ defined in Equation 4. In this way, the inner product of two mapped samples $\phi_{\rm RFF}(\boldsymbol{z}_{1})^{\top}\phi_{\rm RFF}(\boldsymbol{z}_{2})$ provides the approximate Gaussian kernel, so that we can leverage the RFFs mapping $\phi_{\rm RFF}$ to preserve the $\ell_{2}$ distance information through the Gaussian kernel.
Hence, a cosine-Gaussian kernel is adopted for OoD detection, as the Gaussian kernel $k_{\rm gau}$ (or $\phi_{\rm RFF}$) is imposed on top of the cosine kernel $k_{\rm cos}$ (or $\phi_{\rm cos}$), further exploiting the $\ell_{2}$ distance relationships beyond the $\phi_{\rm cos}$-space for OoD detection. As we work with the explicit feature mapping, the non-linearity to $\boldsymbol{z}$ is achieved by $\Phi(\cdot)\triangleq\phi_{\rm RFF}(\phi_{\rm cos}(\cdot))$. With mappings $\phi_{\rm RFF}(\phi_{\rm cos}(\boldsymbol{z}))$, PCA is then executed to compute the reconstruction errors for OoD detection. This detection scheme is dubbed as CoRP (Cosine and RFFs mappings followed by PCA). Algorithm 1 illustrates the complete procedure of the proposed CoP and CoRP for OoD detection. Alternative choices for more kernels are exploited in Section C.2.
To warp up, we devise two effective feature mappings induced from a cosine kernel and a cosine-Gaussian kernel to promote the separability of InD data and OoD data in non-linear feature spaces, inspired by effectiveness of the $\ell_{2}$ normalization and the $\ell_{2}$ distance from a kernel perspective on the KNN detector [7]. Our proposed two feature mappings well characterize the non-linearity in penultimate features $\boldsymbol{z}$ of DNNs between InD and OoD data, enabling PCA to extract an informative subspace w.r.t the mapped features through principal components and the reconstruction errors.
@@ 4.2 Cosine-Gaussian kernel
8:  Calculating the covariance matrix in the mapped $\Phi(\boldsymbol{z})$-space: $\boldsymbol{\Sigma}^{\Phi}=\sum_{i=1}^{N_{\rm tr}}({\color[rgb]{1,0,0}{\Phi(\boldsymbol{z}_{i})}}-\boldsymbol{\mu}^{\Phi})({\color[rgb]{1,0,0}{\Phi(\boldsymbol{z}_{i})}}-\boldsymbol{\mu}^{\Phi})^{\top}$, $\boldsymbol{\mu}^{\Phi}=\frac{1}{N_{\rm tr}}\sum_{i=1}^{N_{\rm tr}}\color[rgb]{1,0,0}{\Phi(\boldsymbol{z}_{i})}$.
@@
10:  Taking the first $q$ columns of $\boldsymbol{U}^{\Phi}$ w.r.t the top-$q$ largest eigenvalues in $\boldsymbol{\Lambda}^{\Phi}$: $\boldsymbol{U}^{\Phi}_{q}=\boldsymbol{U}^{\Phi}\left[:,:q\right]$.
10:  Dimensionality-reduction matrix $\boldsymbol{U}_{q}^{\Phi}$.
11:  Given a new sample $\boldsymbol{\hat{x}}$ and its features $\boldsymbol{\hat{z}}$.
12:  Calculating the reconstruction error:$e^{\Phi}(\boldsymbol{\hat{x}})=\left\|\boldsymbol{U}^{\Phi}_{q}\boldsymbol{U}_{q}^{\Phi\top}({\color[rgb]{1,0,0}{\Phi(\boldsymbol{\hat{z}})}}-\boldsymbol{\mu}^{\Phi})-({\color[rgb]{1,0,0}{\Phi(\boldsymbol{\hat{z}})}}-\boldsymbol{\mu}^{\Phi})\right\|_{2}$.
12:  Reconstruction error $e^{\Phi}$.
Computation complexity
In our method, given any new sample $\boldsymbol{\hat{x}}$ with the penultimate features $\boldsymbol{\hat{z}}$ in inference, to compute the reconstruction error $e^{\Phi}(\boldsymbol{\hat{x}})$, we only need the feature mapping $\Phi$, the projection matrix $\boldsymbol{U}^{\Phi}_{q}$ and the mean mapped training feature vector $\boldsymbol{\mu}^{\Phi}$. Both $\boldsymbol{U}^{\Phi}_{q}$ and $\boldsymbol{\mu}^{\Phi}$ can be pre-calculated and stored in preparation for inference. Therefore, the entire computation cost of CoP and CoRP comes from the construction of the explicit feature mapping $\Phi$ on new features $\boldsymbol{\hat{z}}$.
  • For CoP, its feature mapping $\phi_{\rm cos}$ is an in-place operation and does not require additional computations. Therefore, the time and memory complexity of CoP is ${\cal O}(1)$.
  • For CoRP, the feature mapping $\phi_{\rm RFF}$ of the Gaussian kernel requires $2M$ samplings for $\boldsymbol{\omega}_{i}$ and $u_{i}$, respectively, and $M$ dot-products and $M$ additions. Accordingly, the time and memory complexity of CoRP is ${\cal O}(M)$.
In contrast, regarding the Equation 5 of the KNN detector, all the training features have to be stored at hand and iterated in inference, which implies a heavy $\mathcal{O}(N_{\rm tr})$ time and memory complexity. The ${\cal O}(1)$/${\cal O}(M)$ of CoP/CoRP significantly outperforms the $\mathcal{O}(N_{\rm tr})$ of KNN ($M\ll N_{\rm tr}$). Detailed empirical comparisons are provided in Section 5.1.
In the following, Section 5 exhibits the SOTA performance of our KPCA detector over multiple prevailing detection methods. Section 6 gives an analytical discussion between our covariance-based KPCA and classic KPCA via kernel functions for OoD detection. Due to space limitation, more in-depth investigations on the kernel properties are left in Appendix C, covering ablation studies (Section C.1), alternative kernels (Section C.2) and sensitivity analysis (Section C.3).
5 Experiments on OoD detection
In experiments, our KPCA-based detectors, CoP and CoRP, are firstly compared with KNN [7] in Section 5.1, and show stronger detection performance and cheaper computation costs. In Section 5.2, CoP and CoRP are further compared with the regularized PCA reconstruction error [8], and achieve SOTA OoD detection performance over various prevailing methods. The source code of this work has been publicly released [^1] https://github.com/fanghenshaometeor/ood-kernel-pca. All the experiments are executed on 1 NVIDIA GeForce RTX 3090 GPU.
Datasets
Experiments are executed on the commonly-used small-scale CIFAR10 [39] and large-scale ImageNet-1K benchmarks [40], following the settings in [7; 8]. For CIFAR10 as InD, OoD data sets include SVHN [41], LSUN [42], iSUN [43], Textures [44] and Places365 [45]. For ImageNet-1K as InD, OoD data sets contain iNaturalist [46], SUN [47], Places [45] and Textures [44].
Metrics
For the evaluation metrics, we employ the commonly-used (i) False Positive Rate of OoD samples with 95% true positive rate of InD samples (FPR), and (ii) Area Under the Receiver Operating Characteristic curve (AUROC). The average FPR95 and AUROC values over the selected multiple OoD data sets are viewed as the final comparison metrics.
@@ 5 Experiments on OoD detection
Table 1: The detection performance of different methods (ResNet50 trained on ImageNet-1K).
method OoD data sets AVERAGE
iNaturalist SUN Places Textures
@@
Standard Training
MSP [15] 54.99 87.74 70.83 80.86 73.99 79.76 68.00 79.61 66.95 81.99
ODIN [16] 47.66 89.66 60.15 84.59 67.89 81.78 50.23 85.62 56.48 85.41
Energy [5] 55.72 89.95 59.26 85.89 64.92 82.86 53.72 85.99 58.41 86.17
GODIN [27] 61.91 85.40 60.83 85.60 63.70 83.81 77.85 73.27 66.07 82.02
Mahala [23] 97.00 52.65 98.50 42.41 98.40 41.79 55.80 85.01 87.43 55.47
KNN [7] 59.00 86.47 68.82 80.72 76.28 75.76 11.77 97.07 53.97 85.01
@@
Supervised Contrastive Learning
MSP [15] 32.18 93.30 60.36 84.21 61.68 83.94 50.62 84.68 51.21 86.53
ODIN [16] 23.48 95.80 50.73 88.43 53.99 87.30 41.88 88.60 42.52 90.03
Energy [5] 23.00 95.94 47.56 88.86 51.59 87.58 39.15 89.05 40.33 90.36
SSD [33] 57.16 87.77 78.23 73.10 81.19 70.97 36.37 88.52 63.24 80.09
KNN [7] 30.18 94.89 48.99 88.63 59.15 84.71 15.55 95.40 38.47 90.91
@@
Table 2: Comparisons on the computation complexity between KNN [7] and our CoRP (ResNet50 on ImageNet-1K). Experiments are executed on the same machine for a fair comparison. The nearest neighbor searching of KNN is implemented via Faiss [48].
method time and memorty complexity time consuming (ms, per sample) storage
KNN $\mathcal{O}(N_{\rm tr})$approx 15.59 $\approx$ 20 GiB
CoP ${\cal O}(1)$approx 0.035 $\approx$ 22 MiB
CoRP $\mathcal{O}(M)$approx 0.086 $\approx$ 29 MiB